OpenStudy (marissalovescats):

Express the complex number in trigonometric form. -2+2rad3 i

OpenStudy (marissalovescats):

That's 

OpenStudy (anonymous):

\(z=a+bi=r\exp(i\theta)\) where \[\begin{cases}a=r\cos\theta\\b=r\sin\theta\end{cases}\] \[\begin{cases}-2=r\cos\theta\\2\sqrt3=r\sin\theta\end{cases}\] \(r^2=(-2)^2+(2\sqrt3)^2~~\Rightarrow~~r=4\) Finding \(\theta\) shouldn't be too demanding.

OpenStudy (phi):

find the inverse tangent of imaginary/real

OpenStudy (marissalovescats):

I always know how to find r, but I can't ever find theta for some reason.

OpenStudy (phi):

divide imag/real to get - sqr(3) find atan(-sqr(3)) if you plot -2, 2sqr(3) you see you are in the 2nd quadrant.

OpenStudy (anonymous):

\[-2=4\cos\theta~~\Rightarrow~~\cos\theta=-\frac{1}{2}~~\Rightarrow~~\theta=\frac{2\pi}{3}\text{ or }\frac{4\pi}{3}\] To check which one, use the sine equation: \[2\sqrt3=4\sin\theta~~\Rightarrow~~\sin\theta=\frac{\sqrt3}{2}~~\Rightarrow~~\theta=\frac{\pi}{3}\text{ or }\frac{2\pi}{3}\]

OpenStudy (marissalovescats):

Ah alright, thanks